OpenStudy (anonymous):

Help please! Find the critical numbers of the function f(t)=6t^(2/3)+t^(5/3) there are two critical numbers.

OpenStudy (campbell_st):

ok... so whats the 1st derivative..? and you should also find the 2nd derivative as well

OpenStudy (anonymous):

The derivative of the function would be \[((5x^(2/3)/(3))+((4)/(x^(1/3))\] how do I find the second derivative? @campbell_st

OpenStudy (campbell_st):

wow... I thought \[f'(t) = 6 \times \frac{2}{3} x^{\frac{-1}{3}} + \frac{5}{3} \times t ^{\frac{2}{3}}\] which is \[f'(t) = \frac{4}{\sqrt[3]{t}} + \frac{5t^{\frac{2}{3}}}{3}\]

OpenStudy (anonymous):

I guess I did that wrong... So how do I get the second derivative?

OpenStudy (campbell_st):

or \[f'(t) = \frac{4}{\sqrt[3]{t}} + \frac{5\sqrt[3]{t^2}}{3}\]

OpenStudy (campbell_st):

well don't worry about the 2nd derivative at the minute, its used to test for a max or min stationary point or point of inflection.

OpenStudy (campbell_st):

so you now need to solve for t by letting f'(t) = 0 and the easy solution is t = 0 so thats 1 critical point. you just need to solve for the other. so start by cubing every term \[0 = \frac{64}{t} + \frac{125 t^2}{27}\] then solve for t

OpenStudy (anonymous):

-12/5?

OpenStudy (campbell_st):

then rewrite the equation \[-\frac{64}{t} = \frac{125t^2}{27}\] so \[t^3 = -\frac{1728}{125}\] just take the cube root of both sides

OpenStudy (campbell_st):

thats correct... so the critical points are when x = -12/5 and x = 0 so not sure if you need the actual ordered pair or just the x values.

OpenStudy (anonymous):

Yay! thank you so much! @campbell_st